OpenStudy (wikiemol):

given the matrix | 1 3 1| M = | 2 0 -1| where det(M) = 0 | 1 1 0| the equation MX = 0 will have solutions other than the trivial solution X = 0. How do I find these other solutions?

OpenStudy (wikiemol):

\[M = \left[\begin{matrix}1 & 3 & 1 \\ 2 & 0 &-1\\ 1 & 1 & 0\end{matrix}\right]\]

OpenStudy (amistre64):

by calculating the eigene values and vectors det(M-LI) forms a cubic equation to solve for L, those are your values

OpenStudy (amistre64):

-(-L-1+L-6L) 1-L 2 1 1-L 2 3 -L 1 3 -L 1 −1 -L 1 -1 (L^2-L^3+2-3) L^2 - L^3 + 2 - 3 +L +1- L+ 6L L^2 - L^3 + 6L= 0 L (- L^2 +L + 6) = 0 ; when L=0 L^2 -L - 6 = 0 ; when L = 3, -2

OpenStudy (wikiemol):

Im not sure what eigene values are. In the problem it also says: "Hint: call U, V and W the three rows of M, and observe that M X = 0 if and only if X is orthogonal to the vectors U, V and W." so do they mean there is a way to solve it geometrically? or maybe they are hinting towards something that I havent learned yet.

OpenStudy (amistre64):

hmm, i think im confusing some stuff :) MX = X has to do with eigene values ....

OpenStudy (amistre64):

1x +2y +1z = 0 3x +0y +1z = 0 1x −1y + 0z = 0 per their hint, you have a system of 3 equation in 3 unknowns

OpenStudy (wikiemol):

Well this is a special case. Apparently, if the determinate of M isn't 0, then MX = 0 will only have one trivial solution X = 0. But since it is 0, the worksheet implies that there are other solutions based on the fact that "U, V and W" will form a parallelepiped with no volume. I am not sure what they mean. Here is the worksheet, I am on Problem 3 C http://ocw.mit.edu/courses/mathematics/18-02-multivariable-calculus-fall-2007/exams/prac1a.pdf

OpenStudy (wikiemol):

Ooooooohhh, thank you that explains it.

OpenStudy (wikiemol):

I am kind of confused though on how the determinant of a square matrix finds the volume of a parallelepiped. Was the definition of a determinant made such that A•(B*C) = the determinant of a matrix with the components of the vectors a b and c? or is this derived elsewhere?

OpenStudy (amistre64):

if we take the matrix as is, and row reduce it we get 1 0 1/3 0 1 1/3 0 0 0 this tells us a nul space; or rather, all the vectors that form MX = 0, by solving x and y in terms of z. does that make sense? x = -z/3 y = -z/3 z = z seems to satisfy it

OpenStudy (wikiemol):

I think that makes sense, thanks a lot.

OpenStudy (amistre64):

youre welcome, and good luck .... linear algebra was kind of a pain when i took it :)

OpenStudy (wikiemol):

Haha thanks.